find the fermi energy of copper by assuming that each copper atom contributes one free electron to the electron gas . the density of copper is 8.94 x 10^3kg/m^3


1
Expert's answer
2022-01-26T10:27:37-0500

"E_F=\frac{h^2}{2m}(\frac{3N}{8\pi V})^{\frac 23}=6.97~eV."